# Driving in the Occupancy World: Vision-Centric 4D Occupancy Forecasting and Planning via World Models for Autonomous Driving

### v2

@@ Abstract
World models envision potential future states based on various ego actions. They embed extensive knowledge about the driving environment, facilitating safe and scalable autonomous driving. Most existing methods primarily focus on either data generation or the pretraining paradigms of world models. Unlike the aforementioned prior works, we propose Drive-OccWorld, which adapts a vision-centric 4D forecasting world model to end-to-end planning for autonomous driving. Specifically, we first introduce a semantic and motion-conditional normalization in the memory module, which accumulates semantic and dynamic information from historical BEV embeddings. These BEV features are then conveyed to the world decoder for future occupancy and flow forecasting, considering both geometry and spatiotemporal modeling. Additionally, we propose injecting flexible action conditions, such as velocity, steering angle, trajectory, and commands, into the world model to enable controllable generation and facilitate a broader range of downstream applications. Furthermore, we explore integrating the generative capabilities of the 4D world model with end-to-end planning, enabling continuous forecasting of future states and the selection of optimal trajectories using an occupancy-based cost function. Extensive experiments on the nuScenes dataset demonstrate that our method can generate plausible and controllable 4D occupancy, opening new avenues for driving world generation and end-to-end planning.
@@ 1 Introduction
We evaluate Drive-OccWorld on the nuScenes (Caesar et al. 2020) dataset in terms of vision-centric occupancy and flow forecasting, as well as trajectory planning. In forecasting the sequential occupancy of movable objects and their 3D backward centripetal flow, Drive-OccWorld outperforms previous methods by 2.0% in $\text{mIoU}_{f}$ and 1.9% in $\text{VPQ}_{f}$. For forecasting the occupancy of both movable and static objects based on the OpenOccupancy benchmark (Wang et al. 2023c), it achieves 1.1% gains in $\tilde{\text{mIoU}}_{f}$. Experiments on trajectory planning also demonstrate that Drive-OccWorld can be utilized for safe motion planning.
@@
  • We propose Drive-OccWorld, a vision-centric world model designed for forecasting 4D occupancy and flow, and we explore the integration of the future forecasting capabilities of world models with end-to-end planning.
  • We design a simple yet efficient semantic- and motion-conditional normalization module for semantic enhancement and motion compensation, which improves forecasting and planning performance.
  • We provide a unified conditioning interface that integrates flexible action conditions into future generations, enhancing the controllability of Drive-OccWorld and facilitating a broader range of downstream applications.
@@ 3.1 Preliminary
Figure 3: A detailed overview of the structure of semantic-conditional normalization.
@@
is designed to enhance historical BEV embeddings by incorporating semantic and dynamic information. For example, consider the BEV embedding $F_{-t}^{bev}\in\mathbb{R}^{h\times w\times c}$ at timestamp ${-}t$, where $h$ and $w$ are the spatial resolutions of the BEV, and $c$ is the channel dimension. We first apply layer normalization without affine transformation. Then, the normalized BEV embedding is transformed into $\tilde{F}_{-t}^{bev}$ using a learned affine transformation, with parameters derived from semantic and motion predictions.
As illustrated in Figure 3, we implement a lightweight prediction head to generate voxel-wise semantic probabilities $\mathcal{S}_{-t}\in\mathbb{R}^{h\times w\times d}$ based on $F_{-t}^{bev}$, where $d$ denotes the height of the voxelized 3D space. The semantic labels are encoded into one-hot embeddings and convolved to produce $(\gamma_{-t}^{s},\beta_{-t}^{s})\in\mathbb{R}^{h\times w\times d\times c^{\prime}}$, which serve as learned scale and shift parameters for affine transformation, with $c^{\prime}$ indicating the embedding dimension. This method efficiently integrates BEV embeddings with 3D semantic information.
In motion-conditional normalization, we compensate for the movements of the ego vehicle and other agents across various timestamps. Specifically, the ego-pose transformation matrix $E_{-t}^{+t}$, which accounts for the ego vehicle’s movements from timestamp $-t$ to $+t$, is flattened and encoded into an embedding processed by MLPs to generate affine transformation parameters. Notably, $E_{-t}^{+t}$ is replaced with predicted trajectories $\triangle\tau_{-t}^{+t}$ to prevent ego-status leakage when integrating Drive-OccWorld with planning. To address the movements of other agents, we utilize modulation parameters derived from 3D flow predictions $\mathcal{F}_{-t}\in\mathbb{R}^{h\times w\times d}$.
@@
$\mathcal{W}_{\mathcal{D}}$ is an auto-regressive transformer that predicts the BEV embeddings $F_{+t}^{bev}$ for the future frame $+t$ based on historical BEV features stored in $\mathcal{W}_{\mathcal{M}}$ and the expected action condition $a_{+t}$.
Specifically, $\mathcal{W}_{\mathcal{D}}$ takes learnable BEV queries as input and performs deformable self-attention, temporal cross-attention with historical embeddings, conditional cross-attention with action conditions, and a feedforward network to generate future BEV embeddings. The conditional layer performs cross-attention between BEV queries and action embeddings, which will be illustrated in the following section, injecting action-controllable information into the forecasting process. After obtaining the next BEV embeddings $F_{+t}^{bev}$, prediction heads utilizing the channel-to-height operation (Yu et al. 2023) to predict semantic occupancy and 3D backward centripetal flow $(\mathcal{S}_{+t},\mathcal{F}_{+t})\in\mathbb{R}^{h\times w\times d}$.
@@ 3.3 Action-Controllable Generation
is designed to incorporate heterogeneous action conditions into a coherent embedding, inspired by (Gao et al. 2024; Wang et al. 2024b). We first encode the required actions into Fourier embeddings (Tancik et al. 2020), which are concatenated and fused through additional learned projections to align with the dimensions of the conditional cross-attention layers in $\mathcal{W}_{\mathcal{D}}$. This method efficiently integrates flexible conditions into controllable generation.
@@ 3.4 End-to-End Planning with World Model
Existing world models primarily focus on either data generation or the pertaining paradigms for autonomous driving. Although a recent pioneering work, Drive-WM (Wang et al. 2024b), proposed integrating generated driving videos with an image-based reward function to plan trajectories, the geometric 3D features of the environment are not fully exploited for motion planning. Given the capabilities of future occupancy forecasting provided by our world model, as illustrated in Figure 2, we introduce an occupancy-based planner to sample the occupied grids of agents and drivable areas for safety constraints. Additionally, future BEV embeddings are used to learn a cost volume that accounts for fine-grained 3D structure, providing more comprehensive environmental information for safe planning.
@@
is designed to ensure the safe driving of the ego vehicle. It consists of multiple cost factors: (1) Agent-Safety Cost constrains the ego vehicle from colliding with other agents, such as pedestrians and vehicles. It penalizes trajectory candidates that overlap with grids occupied by other road users. Additionally, trajectories that are too close to other agents, in terms of lateral or longitudinal distance, are also restricted to avoid potential collisions. (2) Road-Safety Cost ensures the vehicle remains on the road. It extracts road layouts from occupancy predictions, penalizing trajectories that fall outside the drivable area. (3) Learned-Volume Cost is inspired by ST-P3 (Hu et al. 2022). It utilizes a learnable head based on $F_{+t}^{bev}$ to generate a cost volume, providing a more comprehensive evaluation of the complex world.
The total cost function is the summation of the above cost factors. Similar to ST-P3, a sampler generates a set of trajectory candidates $\tau_{+t}^{*}$ based on high-level commands, and then the planner $\mathcal{P}$ selects the optimal trajectory $\tau_{+t}$ with the minimum cost, considering both agent and road safety.
@@
is introduced to further refine the trajectory using BEV embeddings $F_{+t}^{bev}$ in the latent space. We encode $\tau_{+t}$ into an embedding and concatenate it with a command embedding to form an ego query, which performs cross-attention with $F_{+t}^{bev}$ to extract fine-grained representations of the environment. The final trajectory is predicted based on the refined ego query through MLPs.
The planning loss $\mathcal{L}_{plan}$ consists of three components: a max-margin loss introduced by (Sadat et al. 2020) to constrain the safety of trajectory candidates $\tau_{+t}^{*}$, a naive $l_{2}$ loss for imitation learning, and a collision loss that ensures the planned trajectory avoids grids occupied by obstacles.
@@ 4.1 Setup
We conduct experiments on the nuScenes (Caesar et al. 2020) and nuScenes-Occupancy (Wang et al. 2023c) datasets. More details regarding the experimental implementation are provided in the appendix.
@@
We validate the effectiveness of Drive-OccWorld on 4D occupancy and flow forecasting, encompassing both inflated and fine-grained occupancy formats, as well as trajectory planning. (1) Inflated Occupancy and Flow Forecasting is introduced in Cam4DOcc (Ma et al. 2024b), predicting the future states of movable objects with dilated occupancy patterns. The dataset is constructed by transforming movable objects at different timestamps into the present coordinate system, voxelizing the 3D space, and attaching semantic and instance labels to the grids using bounding box annotations. The 3D backward centripetal flow indicates the direction of the voxels in the current frame toward their corresponding 3D instance centers within the previous timestamp. (2) Fine-grained Occupancy Forecasting utilizes voxel-level semantic occupancy annotations provided by nuScenes-Occupancy, which include both movable objects and static environments. (3) End-to-end Planning follows the open-loop evaluation on nuScenes.
@@
Method $\text{mIoU}_{c}$ $\text{mIoU}_{f}$ (2 s) $\tilde{\text{mIoU}}_{f}$ $\text{VPQ}_{f}$ $\text{VPQ}_{f}^{*}$
SPC 1.3 failed failed
OpenOccupancy-C (Wang et al. 2023c) 12.2 11.5 11.7
PowerBEV-3D (Li et al. 2023a) 23.1 21.3 21.9
Cam4DOcc (Ma et al. 2024b) 31.3 26.8 28.0 21.4
Drive-OccWorldA (Ours) 29.4 28.6 28.7 22.6 32.0
Drive-OccWorldP (Ours) 29.6 28.8 29.0 23.3 33.2
SPC: SurroundDepth (Wei et al. 2023a) + PCPNet (Luo et al. 2023)
+ Cylinder3D (Zhu et al. 2021)
Table 1: Comparisons of Inflated Occupancy and Flow Forecasting on the nuScenes validation set, with the top two results highlighted in bold and underlined.
@@
DriveOccWorldA (Ours) 12.3 16.8 14.5 10.2 16.7 13.5 10.4
DriveOccWorldP (Ours) 12.6 17.5 15.1 10.3 17.9 14.1 10.8
@@
Table 2: Comparisons of Fine-grained Occupancy Forecasting on nuScenes-Occupancy, with the top two results highlighted in bold and underlined. GMO indicates general movable objects, while GSO refers to general static objects.
@@ 4.2 Main Results of 4D Occupancy Forecasting
Inflated Occupancy and Flow Forecasting.
Table 1 presents comparisons of inflated occupancy and flow forecasting on the nuScenes dataset. Although Drive-OccWorld performs slightly weaker on the current moment’s result, $\text{mIoU}_{c}$, it outperforms Cam4DOcc on $\text{mIoU}_{f}$ by 2.0%, demonstrating a stronger ability to forecast future states.
For future flow predictions, Drive-OccWorldP continuously outperforms previous methods by 1.9% on $\text{VPQ}_{f}$, indicating a superior capability for modeling the motions of dynamic objects. Furthermore, by integrating the center clustering technique, Drive-OccWorld achieves remarkable results on $\text{VPQ}_{f}^{*}$, showcasing the effectiveness of instance proposals association through center clustering.
@@
trajectory velocity angle command
@@
Fine-grained Occupancy Forecasting.
Table 2 presents comparisons of fine-grained occupancy forecasting on the nuScenes-Occupancy. The results demonstrate that Drive-OccWorld achieves the best performance compared to all other approaches. Notably, Drive-OccWorldP outperforms Cam4DOcc by 1.6% and 1.1% on mIoU for general movable objects at current and future timestamps, respectively, illustrating its ability to accurately locate movable objects for safe planning. Figure 4 provides qualitative results of the occupancy forecasting and flow predictions across frames.
@@
In Table 3, we examine controllability under diverse action conditions. Compared to the baseline variant, injecting any action condition yields gains. It is worth noting that low-level conditions, i.e., trajectory and velocity, provide greater improvements in the result $\tilde{\text{mIoU}}_{f}$ for future forecasting. In contrast, the highest-level command condition improves $\text{mIoU}_{c}$ results at the current timestamp but offers limited enhancement for future predictions. We explain that incorporating more low-level conditions, such as trajectory, can provide more concrete actions for the ego vehicle to understand its interactions with the world, thereby effectively enhancing future forecasting.
@@
Table 4: Planning upper bound of the Drive-OccWorld when using GT trajectory as action condition.
Interestingly, as shown in Table 4, using ground-truth trajectories as action conditions leads to superior planning results compared to using predicted trajectories. Conversely, the opposite trend is observed in occupancy and flow forecasting quality. Comparisons of line 2 and line 7 in Table 3 reveal that utilizing predicted trajectories instead of ground-truth ones yields slightly greater improvements in forecasting quality. This observation is further supported by the results in Tables 1 and 2, where Drive-OccWorldP outperforms Drive-OccWorldA. We suggest that the performance gains in occupancy and flow quality may result from the trajectory constraints imposed on the BEV features when using predicted trajectories. This finding indicates that applying trajectory prediction enhances perception performance as well, consistent with the results in UniAD.

### v3

@@ Abstract
World models envision potential future states based on various ego actions. They embed extensive knowledge about the driving environment, facilitating safe and scalable autonomous driving. Most existing methods primarily focus on either data generation or the pretraining paradigms of world models. Unlike the aforementioned prior works, we propose Drive-OccWorld, which adapts a vision-centric 4D forecasting world model to end-to-end planning for autonomous driving. Specifically, we first introduce a semantic and motion-conditional normalization in the memory module, which accumulates semantic and dynamic information from historical BEV embeddings. These BEV features are then conveyed to the world decoder for future occupancy and flow forecasting, considering both geometry and spatiotemporal modeling. Additionally, we propose injecting flexible action conditions, such as velocity, steering angle, trajectory, and commands, into the world model to enable controllable generation and facilitate a broader range of downstream applications. Furthermore, we explore integrating the generative capabilities of the 4D world model with end-to-end planning, enabling continuous forecasting of future states and the selection of optimal trajectories using an occupancy-based cost function. Comprehensive experiments conducted on the nuScenes, nuScenes-Occupancy, and Lyft-Level5 datasets illustrate that our method can generate plausible and controllable 4D occupancy, paving the way for advancements in driving world generation and end-to-end planning.
@@ 1 Introduction
We evaluate Drive-OccWorld using the nuScenes, nuScenes-Occupancy, and Lyft-Level5 datasets. It outperforms previous methods by 9.5% in $\text{mIoU}_{f}$ and 5.1% in $\text{VPQ}_{f}$ on nuScenes, by 6.1% in $\text{mIoU}_{f}$ and 5.2% in $\text{VPQ}_{f}$ on Lyft-Level5 for forecasting the inflated occupancy of movable objects and their 3D backward centripetal flow. On the nuScenes-Occupancy benchmark, it achieves a 4.3% improvement in fine-grained occupancy forecasting. Additionally, experiments on trajectory planning indicate that Drive-OccWorld is effective for safe motion planning.
@@
  • We propose Drive-OccWorld, a vision-centric world model designed for forecasting 4D occupancy and dynamic flow, achieving new state-of-the-art performance on both the nuScenes and Lyft benchmarks.
  • We develop a simple yet efficient semantic- and motion-conditional normalization module for semantic enhancement and motion compensation, which improves forecasting and planning performance.
  • We incorporate flexible action conditions into Drive-OccWorld to enable action-controllable generation and explore integrating the world model with an occupancy-based planner for continuous forecasting and planning.
@@ 3.1 Preliminary
Figure 3: Overview of semantic-conditional normalization.
@@
is designed to enhance historical BEV embeddings by incorporating semantic and dynamic information. For example, consider the BEV embedding $\bm{F}^{bev}\in\mathbb{R}^{h\times w\times c}$, where $h$ and $w$ are the spatial resolutions of the BEV, and $c$ denotes the channel dimension. We first apply layer normalization without affine mapping, then modulate it into $\tilde{\bm{F}}^{bev}$ using an adaptive affine transformation, with the scale and shift parameters $(\gamma^{*},\beta^{*})$ derived from semantic or motion labels:
$\tilde{\bm{F}}^{bev}=\gamma^{*}\cdot\textit{LayerNorm}(\bm{F}^{bev})+\beta^{*}$ (5)
Specifically, for semantic-conditional normalization, $(\gamma^{s},\beta^{s})$ are inferred from voxel-wise semantic predictions. As illustrated in Figure 3, we implement a lightweight head along with the argmax function to predict voxel-wise semantic labels $\mathcal{S}\in\mathbb{R}^{h\times w\times d\times 1}$, where $d$ denotes the height of the voxelized 3D space. The semantic labels are encoded as one-hot embeddings and convolved to produce modulation parameters for the affine transformation as Eq. 5. This method efficiently enhances the semantic discrimination of BEV embeddings, as demonstrated in the experiments.
In motion-conditional normalization, we account for the movements of both the ego vehicle and other agents across various timestamps. Specifically, the ego-pose transformation matrix $E_{-t}^{+t}=[R_{-t}^{+t},T_{-t}^{+t}]$, which represents the rotation and translation of the ego vehicle from timestamp $-t$ to $+t$, is flattened and encoded into an embedding processed by MLPs to generate affine transformation parameters $(\gamma^{e},\beta^{e})$. To address the movements of other agents, we predict voxel-wise 3D backward centripetal flow $\mathcal{F}\in\mathbb{R}^{h\times w\times d\times 3}$ that points from the voxel at time $t$ to its corresponding 3D instance center at $t-1$, and encode it into $(\gamma^{f},\beta^{f})$ for fine-grained motion-aware normalization using Eq. 5.
@@
$\mathcal{W}_{\mathcal{D}}$ is an auto-regressive transformer that predicts the BEV embeddings $\bm{F}_{+t}^{bev}$ for the future frame $+t$ based on historical BEV features stored in $\mathcal{W}_{\mathcal{M}}$ and the expected action condition $a_{+t}$.
Specifically, $\mathcal{W}_{\mathcal{D}}$ takes learnable BEV queries as input and performs deformable self-attention, temporal cross-attention with historical embeddings, conditional cross-attention with action conditions, and a feedforward network to generate future BEV embeddings. The conditional layer performs cross-attention between BEV queries and action embeddings, which will be illustrated in the following section, injecting action-controllable information into the forecasting process. After obtaining the next BEV embeddings $\bm{F}_{+t}^{bev}$, prediction heads utilizing the channel-to-height operation (Yu et al. 2023) to predict semantic occupancy and 3D backward centripetal flow $(\mathcal{S}_{+t},\mathcal{F}_{+t})\in\mathbb{R}^{h\times w\times d}$.
@@ 3.3 Action-Controllable Generation
is designed to incorporate heterogeneous action conditions into a coherent embedding, inspired by (Gao et al. 2024; Wang et al. 2024b). We first encode the required actions via Fourier embeddings (Tancik et al. 2020), which are then concatenated and fused via learned projections to align with the dimensions of the conditional cross-attention layers in $\mathcal{W}_{\mathcal{D}}$. This method enables efficient integration of flexible action conditions into controllable generation, with experiments demonstrating that the unified interface with conditional cross-attention provides superior controllability compared to other approaches such as additive embeddings.
@@ 3.4 End-to-End Planning with World Model
Existing world models primarily focus on either data generation or the pertaining paradigms for autonomous driving. Although a recent pioneering work, Drive-WM (Wang et al. 2024b), proposed integrating generated driving videos with an image-based reward function to plan trajectories, the geometric 3D features of the environment are not fully exploited for motion planning. Leveraging the future occupancy forecasting capabilities of our world model, as illustrated in Figure 2, we introduce an occupancy-based planner that samples occupied grids of agents and drivable areas to enforce safety constraints. Additionally, we employ a learned-volume cost to provide a more comprehensive evaluation of the environment for safe planning.
@@
is designed to ensure the safe driving of the ego vehicle. It consists of multiple cost factors: (1) Agent-Safety Cost constrains the ego vehicle from colliding with other agents, such as pedestrians and vehicles. It penalizes trajectory candidates that overlap with grids occupied by other road users. Additionally, trajectories that are too close to other agents, in terms of lateral or longitudinal distance, are also restricted to avoid potential collisions. (2) Road-Safety Cost ensures the vehicle remains on the road. It extracts road layouts from occupancy predictions, penalizing trajectories that fall outside the drivable area. (3) Learned-Volume Cost is inspired by ST-P3 (Hu et al. 2022). It employs a learnable head based on $\bm{F}_{+t}^{bev}$ to generate a 2D cost map, enabling a more comprehensive evaluation of occupancy grids in complex environments.
The total cost function is the summation of the above cost factors. Following the approach of ST-P3, a trajectory sampler generates a set of candidate trajectories $\tau_{+t}^{*}\in\mathbb{R}^{N_{\tau}\times 2}$ distributed across the 2D grid map surrounding the ego vehicle, guided by high-level commands. Subsequently, the trajectory planner $\mathcal{P}$ selects the optimal trajectory $\tau_{+t}$ by minimizing the total cost function, while simultaneously ensuring agent and road safety.
@@
is introduced to further refine the trajectory using the latent features of BEV embeddings $\bm{F}_{+t}^{bev}$. We encode $\tau_{+t}$ into an embedding and concatenate it with a command embedding to form an ego query, which performs cross-attention with $\bm{F}_{+t}^{bev}$ to extract fine-grained representations of the environment. The final trajectory is predicted based on the refined ego query through MLPs.
The planning loss $\mathcal{L}_{plan}$ consists of three components: a max-margin loss introduced by (Sadat et al. 2020) to constrain the safety of trajectory candidates $\tau_{+t}^{*}$, a naive $l_{2}$ loss for imitation learning, and a collision loss that ensures the planned trajectory avoids grids occupied by obstacles. Notably, when performing end-to-end planning, we utilize the predicted trajectories as action conditions for both training and testing. This approach not only prevents GT ego actions from leaking into the planner but also facilitates model learning using predicted trajectories, enabling improved performance during testing.
Method Inflated GMO Fine-Grained GMO
nuScenes Lyft-Level5 nuScenes-Occupancy
$\text{mIoU}_{c}$ $\text{mIoU}_{f}$ (2 s) $\tilde{\text{mIoU}}_{f}$ $\text{VPQ}_{f}$ $\text{IoU}_{c}$ $\text{mIoU}_{f}$ (0.8 s) $\tilde{\text{mIoU}}_{f}$ $\text{VPQ}_{f}$ $\text{mIoU}_{c}$ $\text{mIoU}_{f}$ (2 s) $\tilde{\text{mIoU}}_{f}$
SPC 1.3 failed failed 1.4 failed failed 5.9 1.1 1.1
OpenOccupancy-C (Wang et al. 2023c) 12.2 11.5 11.7 14.0 13.5 13.7 10.8 8.0 8.5
PowerBEV-3D (Li et al. 2023a) 23.1 21.3 21.9 20.0 26.2 24.5 25.1 27.4 5.9 5.3 5.5
Cam4DOcc (Ma et al. 2024b) 31.3 26.8 28.0 18.6 36.4 33.6 34.6 28.2 11.5 9.7 10.1
Drive-OccWorldA (Ours) 39.7 36.3 37.3 23.7 40.6 39.3 40.0 32.2 13.6 11.9 12.3
Drive-OccWorldP (Ours) 39.8 36.3 37.4 25.1 40.9 39.7 40.6 33.4 13.6 12.0 12.4
SPC: SurroundDepth (Wei et al. 2023a) + PCPNet (Luo et al. 2023) + Cylinder3D (Zhu et al. 2021)
Table 1: Comparisons of Inflated GMO and Flow Forecasting on the nuScenes and Lyft-Level5 datasets, and Fine-Grained GMO Forecasting on the nuScenes-Occupancy dataset, with the top two results highlighted in bold and underlined text.
@@ 4.1 Setup
We conduct experiments on the nuScenes (Caesar et al. 2020), nuScenes-Occupancy (Wang et al. 2023c), and Lyft-Level5 datasets. More details regarding the experimental implementation are provided in the appendix.
@@
We validate Drive-OccWorld’s effectiveness in 4D occupancy and flow forecasting, which includes inflated and fine-grained occupancy formats, as well as trajectory planning. (1) Inflated GMO and Flow Forecasting is presented in Cam4DOcc (Ma et al. 2024b), predicting the future states of general movable objects (GMO) with dilated occupancy patterns. The dataset is constructed by transforming movable objects at different timestamps into the present coordinate system, voxelizing the 3D space, and labeling grids with semantic and instance annotations. Voxel grids within bounding boxes are marked as GMO, and the 3D backward centripetal flow indicates the direction of voxels in the current frame toward their corresponding 3D instance centers from the previous timestamp. (2) Fine-Grained GMO Forecasting predicts occupied grids of GMO using voxel-wise labels from the nuScenes-Occupancy dataset. (3) Fine-grained GMO and GSO Forecasting predicts the voxel-wise labels for both general movable objects (GMO) and general static objects (GSO) based on fine-grained annotations from nuScenes-Occupancy. (4) End-to-end Planning follows the open-loop evaluation on nuScenes.
@@
DriveOccWorldA (Ours) 16.6 20.2 18.4 14.3 21.4 17.8 14.9
DriveOccWorldP (Ours) 16.9 20.2 18.5 14.3 21.2 17.8 14.9
@@
Table 2: Comparisons of Fine-Grained GMO and GSO Forecasting on nuScenes-Occupancy dataset.
@@ 4.2 Main Results of 4D Occupancy Forecasting
Inflated GMO and Flow Forecasting.
Table 1 presents comparisons of inflated GMO and flow forecasting on the nuScenes and Lyft-Level5 datasets. Drive-OccWorld outperforms previous methods on various time intervals, including the performance on the current moment and future timestamps. For instance, it surpasses Cam4DOcc on the $\tilde{\text{mIoU}}_{f}$ metric by 9.4% and 6% on nuScenes and Lyft-Level5, respectively. The results demonstrate Drive-OccWorld’s superior ability to forecast future world states.
For future flow predictions, Drive-OccWorldP outperforms previous SoTA methods by 5.1% and 5.2% on $\text{VPQ}_{f}$ for nuScenes and Lyft-Level5, respectively, indicating superior capability in modeling dynamic object motions.
Fine-Grained GMO Forecasting.
Fine-grained GMO forecasting poses more challenges than inflated GMO forecasting, since it requires predicting more fine-grained voxel labels than the bounding-box level labels. Table 1 presents comparisons of fine-grained GMO forecasting on the nuScenes-Occupancy dataset. The results show that Drive-OccWorld still outperforms previous methods across all time intervals, showcasing Drive-OccWorld’s ability to forecast more fine-grained world states.
@@
traj vel angle cmd
@@
Fine-grained GMO and GSO Forecasting.
Table 2 presents comparisons of fine-grained GMO and GSO forecasting on the nuScenes-Occupancy dataset. The results demonstrate that Drive-OccWorld achieves the best performance among all approaches. Notably, Drive-OccWorldP outperforms Cam4DOcc by 5.9% and 5.1% on mIoU for general movable objects (GMO) at current and future timestamps, respectively, illustrating its ability to accurately locate movable objects for safe navigation. Figure 4 provides qualitative results of fine-grained occupancy forecasting and flow predictions across frames.
@@
Table 3 examines controllability under various action conditions. Injecting any action condition improves results compared to the baseline. Low-level conditions, such as trajectory and velocity, significantly enhance future forecasting ($\tilde{\text{mIoU}}_{f}$), while high-level conditions, such as commands, boost results at the current moment ($\text{mIoU}_{c}$) but have limited impact on future predictions. We explain that incorporating more low-level conditions helps the world model better understand how the ego vehicle interacts with the environment, thereby improving forecasting performance.
@@
Table 4: Planning upper bound when using GT trajectory.
Table 4 shows that using ground-truth trajectories as action conditions yields better planning results than predicted trajectories. However, using predicted trajectories slightly improves occupancy and flow forecasting quality, as indicated by comparisons in Table 3 (line 2 vs. line 7) and supported by Tables 1 and 2, where Drive-OccWorldP outperforms Drive-OccWorldA. This performance gain may stem from planning constraints associated with predicted trajectories, allowing the planner to perform cross-attention between trajectories and BEV features. This process constrains the BEV features to account for ego motion, thereby enhancing perception performance. Additionally, using predicted trajectories during training improves model learning to boost performance during testing.